# Entanglement engineering of optomechanical systems by reinforcement learning

### v1

@@ Abstract
Entanglement is fundamental to quantum information science and technology, yet controlling and manipulating entanglement — so-called entanglement engineering — for arbitrary quantum systems remains a formidable challenge. There are two difficulties: the fragility of quantum entanglement and its experimental characterization. We develop a deep reinforcement-learning (RL) approach to entanglement engineering, in which feedback control together with weak continuous measurement and partial state observation is exploited to generate and maintain desired entanglement. We employ quantum optomechanical systems with linear or nonlinear photon-phonon interactions to demonstrate the workings of our machine-learning-based entanglement engineering protocol. In particular, the RL agent sequentially interacts with one or multiple parallel quantum optomechanical environments, collects trajectories, and updates the policy to maximize the accumulated reward to create and stabilize quantum entanglement over an arbitrary amount of time. The machine-learning-based control principle is applicable to entanglement engineering of experimental quantum systems in general.
I Introduction
@@
Entanglement [1, 2, 3, 4, 5] is fundamental to all fields in quantum information science such as quantum sensing [6], quantum computation [7], and quantum networks [8, 9, 10, 11, 12, 13]. However, the inherent fragility of quantum entanglement and coherence [14] poses significant challenges for experimental applications. For example, in quantum computing, depending on the applications, the quantum states of the quantum gates need last for a finite amount of time [15, 16, 17, 18, 19, 20], making critical maintaining the entanglement after its creation. Moreover, the transition from noisy intermediate-scale systems [21] to large-scale, fault-tolerant systems [16] requires sophisticated entanglement engineering strategies to establish and maintain entanglement in the presence of noise and decoherence.
In this paper, we propose a reinforcement learning (RL) based approach to entanglement engineering, with the goal of developing a feedback control strategy to create and maintain quantum entanglement. In the next three paragraphs, we briefly describe the background of the three pertinent topics: quantum control, deep RL, and quantum measurement.
Quantum control.
Quantum control [22] is essential to quantum engineering and technology [23, 24, 25], where open-loop control [26] has been successfully demonstrated with methods such as gradient-ascent pulse engineering (GRAPE) [27] in spin systems [28], coupled qubits [29], Jaynes-Cummings systems [30], and qubit-cavity lattices [31]. Recently, the open-loop GRAPE algorithm has been extended to feedback GRAPE [32] based on gradient ascent of quantum dynamics for state engineering under strongly stochastic measurement. Open-loop control, however, requires a differentiable model of the quantum dynamics that may not always be available. In realistic situations where such a model is not available, closed-loop feedback control strategies conditioned on experimental measurement outcomes can be applied. Combined with data-driven machine learning, feedback control has been implemented in experiments in a model-free fashion [33, 34, 35].
Deep reinforcement learning.
In general, RL is a machine-learning paradigm based on a trial-and-error learning process, incorporating traditional optimal control to maximize the accumulated reward. The use of deep neural networks in the learning process leads to deep RL, which explores and exploits the available measurement data to search for a globally optimal policy. In deep RL, many algorithms are available such as deep-Q network (DQN) [33], deep deterministic policy gradient (DDPG) [36], and trust region proximal optimization (TRPO) [37]. A state-of-the-art deep RL algorithm for continuous control is proximal policy optimization (PPO) [38], whose performance can exceed that of TRPO. Incorporating recurrent neural networks [39] into the PPO algorithm leads to improved performance [40]. In recent years, measurement-based feedback control with deep RL has been applied to quantum systems for tasks such as quantum error correction for discrete gates [41], state preparation and stabilization for a single particle [42, 43, 44, 45] with an unstable potential [42] or a double-well potential [43], discrimination between entangled states [46] for quantum meteorology, and long-distance entanglement distribution on quantum networks [47]. Experimentally, time scales of the RL action sequences shorter than the coherence time of the underlying quantum system have been realized, rendering feasible real-time deep-RL feedback control [48].
Quantum measurement.
In quantum systems, projective measurement can be used to extract the full information about the quantum state but, as a back action, the quantum state will collapse after the measurement [49]. To avoid a complete collapse, one can exploit weak measurements [50, 51], in which the probe is weakly coupled to the system to yield partial information about the quantum state. Examples of weak measurements include continuous monitoring [52] of driven dissipative quantum-optical systems - a basic component of quantum meteorology [53, 54]. A form of weak measurement, the so-called weak continuous measurement (WCM), is fundamental to a broad range of applications. For example, WCM has been used to detect the quadrature operators [55], Wigner [55] and Husimi Q functions [56] with a homodyne apparatus [57], rendering observing both pure [58] and mixed [50] quantum states experimentally feasible. WCM has been experimentally implemented by a weak-field homodyne detector [59, 55, 57] to measure the photon-number statistical distribution over the Fock basis. In another example, WCM has been realized in an atomic spin ensemble [51] via Faraday rotation of an off-resonance probe beam to create and probe nonclassical spin state and dynamics. The concept of WCM has also been used to develop fundamental theories, such as Heisenberg’s measurement-disturbance relationship [60] and error-disturbance uncertainty relation [61]. Because of the typical time scales of the quantum dynamics, WCM cannot be regarded as occurring instantaneously [62]. Theoretically, the impact of WCM on the underlying quantum system can be described by the stochastic master equation [62].
Our goal is to create and stabilize quantum systems at predetermined entanglement levels, effectively realizing controllable entanglement. To address this foundational problem in entanglement engineering, we develop a model-free, deep RL-based feedback control framework to create and stabilize entanglement based on partial observation in quantum optomechanical systems, where a stochastic master equation, subject to WCM, governs the underlying quantum system. We consider both linear and nonlinear interactions between the optical and mechanical modes, where the observation is WCM photocurrent for the former and the expectation value of the photon number for the latter. We demonstrate that, under the actions of the PPO or recurrent PPO RL agent, entanglement between the quantum optical and mechanical modes can be created and maintained.
@@
More specifically, under the strong laser approximation, the interaction resulting from the radiation pressure between the cavity and the mechanical oscillator modes can be linearized and described by the beam-splitter Hamiltonian. During the training phase, the PPO agent interacts with parallel quantum environments and collects the subsequent data by episodic learning, with the observation being the WCM current. The deep-RL method can extract useful information from the measurement current, which is encoded in the Wiener process, and achieve the target entanglement engineering in a model-free manner for the quantum system that is dissipative due to coupling to the vacuum bath and is driven by a laser. In the testing phase, with the agent interacting and observing a single quantum environment, we demonstrate that the entanglement-engineering performance of our deep-RL method with WCM observation greatly exceeds that of both state-based Bayesian methods [63, 43] and random control.
When the driving laser field is not strong, the quantum optomechanical interaction is nonlinear [64, 65]. In this case, we articulate two training phases for nonlinear entanglement engineering. The first is model-based and infers the entanglement, dubbed the target-generating phase, where the observation of the PPO agent is the logarithmic negativity and the reward function is constructed to limit the high-level excitation and facilitate entanglement learning. (Note that direct experimental measurement of the logarithmic negativity is currently not available.) The time series of the expected photon number in the regime from converged training episodes is selected as the target for the next phase. The second phase is then the target-utilization phase, where the recurrent PPO observes the expected photon number and obtains the reward only based on the target expected photon number obtained from the last phase. In this framework, the recurrent PPO controls the quantum state in the low-energy regime with the desired entanglement stabilized.
In Sec. II, we describe the noise model of the quantum optomechanical system to generate training data learned by the RL agent. In Sec. III, we introduce our deep RL control framework for quantum optomechanical systems with linear interactions. In Sec. IV, we present the performance results of deep RL control for optomechanical systems with nonlinear interactions. Pertinent issues such as benchmarks, the compatibility of long time horizon, effects of disturbance, and mixture tolerance for the initial state are addressed. In Sec. VI, conclusions and an extensive discussion about limitations and challenges are presented.
II Quantum optomechanical systems and dynamics
II.1 Hamiltonian
We consider a cavity optomechanical system that consists of a single-mode cavity and a movable end mirror, as shown in Fig 1 (a). The optical cavity h as the frequency $\omega_{c}$ and the optical field exerts a radiation pressure on the mirror. The cavity mode is driven externally by a coherent laser field with frequency $\omega_{L}$. The mirror’s quantized center-of-mass motion is described by a harmonic oscillator of frequency $\omega_{m}$. In the rotating frame of the laser, the Hamiltonian describing the coupling between the optical cavity and mechanical oscillator modes is given by [64, 65]
$\tilde{H}_{nl}=$ $-\hbar\Delta\hat{a}^{\dagger}\hat{a}+\hbar\omega_{m}\hat{b}^{\dagger}\hat{b}+\hbar g_{0}(\hat{b}^{\dagger}+\hat{b})\hat{a}^{\dagger}\hat{a}$
$+\hbar\alpha_{L}(\hat{a}^{\dagger}+\hat{a}),$ (1)
where $\hat{a}$ and $\hat{b}$ are the annihilation operators of the cavity and mechanical mode, respectively, $\hat{a}^{\dagger}$ and $\hat{b}^{\dagger}$ are the corresponding creation operators. The frequency detuning of the cavity is $\Delta\equiv\omega_{L}-\omega_{c}$. The nonlinear coupling $g_{0}$ arises from the radiation pressure force between the light and the movable mirror (details given in Appendix A), and $\alpha_{L}$ is the real amplitude of the driven electromagnetic field. We set $g_{0}>\kappa$ so that the single-photon optomechanical coupling rate $g_{0}$ exceeds the coupling strength $\kappa$ between the cavity and the vacuum bath. This condition guarantees observable nonlinear quantum effects [66]. In the rotating frame, the displacement transformation $\hat{a}\approx\bar{\alpha}_{c}+\delta\hat{a}$ can be applied, where $|\bar{\alpha}_{c}|\gg 1$ is the amplitude of the light field inside the cavity induced by the strong laser, and $\delta\hat{a}$ denotes the excitation or the shifted oscillator on top of the large coherent state with the amplitude $\bar{\alpha}_{c}$. If the third- and higher-order terms can be neglected, the nonlinear coupling can be reduced to the linear interaction [67, 64], leading to the following linearized Hamiltonian:
$\tilde{H}_{l}\approx-\hbar\Delta\delta\hat{a}^{\dagger}\delta\hat{a}+\hbar\omega_{m}\hat{b}^{\dagger}\hat{b}+\hbar G(\delta\hat{a}^{\dagger}+\delta\hat{a})(\hat{b}^{\dagger}+\hat{b}),$ (2)
where the coefficient $G\equiv g_{0}\bar{\alpha}_{c}$ can be tuned by the amplitude of the incoming laser (a time-dependent modulation) [68]. In the red-detuned regime $\Delta=-\omega_{m}$, all off-resonance terms can be omitted, leading to the following beam-splitter or “swap” Hamiltonian:
$\tilde{H}_{bs}\approx\hbar\omega_{m}\delta\hat{a}^{\dagger}\delta\hat{a}+\hbar\omega_{m}\hat{b}^{\dagger}\hat{b}+\hbar G(\delta\hat{a}^{\dagger}\hat{b}+\hat{b}^{\dagger}\delta\hat{a}),$ (3)
where the interaction term describes the state transfer between photons and phonons in the strong coupling regime for $G>\kappa$, with $\kappa$ being the decay rate of the cavity mode to the vacuum bath at zero temperature.
II.2 Quantum dynamics of an open optomechanical system
An experimental optomechanical system is effectively an open quantum system interacting with the vacuum bath under WCM with the operators $\hat{C}_{n}\equiv\sqrt{\eta}\hat{P}_{n}$ by homodyne detection [69, 45], where $\hat{P}_{n}=|n\rangle\langle n|$ with $n=0,1$ (linear) or $n=0,1,\ldots,9$ (nonlinear) is the measurement operator on the Fock state and $\eta$ denotes the measurement rate. The quantum dynamics of this system are described by the stochastic master equation (SME) [70, 45, 71, 22, 62]:
$d\rho$ $=\frac{1}{i\hbar}[\tilde{H},\rho]dt+\mathcal{L}_{env}\;\rho dt$
$+\sum_{n}\mathcal{D}(\hat{C}_{n})\rho dt+\sum_{n}\mathcal{H}(\hat{C}_{n})\rho{\color[rgb]{0,0,1}dW_{n}},$ (4)
where the Hamiltonian is $\tilde{H}=\tilde{H}_{bs}$ or $\tilde{H}_{nl}$ and $\rho$ is a density operator in the Hilbert space. Under the Born-Markov approximation [72, 73], which requires the system-bath coupling to be weak and the correlation time of the bath to be much shorter than a characteristic timescale of system-bath interactions, the Markovian master equation, i.e., the first two terms in Eq. (II.2), has the Lindblad form [72]. At absolute zero temperature, the following environmental operator $\mathcal{L}_{env}\;\rho$ can be introduced to describe the coupling between the system and vacuum bath:
$\mathcal{L}_{env}\;\rho=\kappa\mathcal{D}(\hat{a})\rho+\gamma\mathcal{D}(\hat{b})\rho,$
where the cavity and oscillator modes are coupled to the vacuum bath with the strength $\kappa$ and $\gamma$, respectively [65]. Note that the WCM process described by the last two terms of Eq. (II.2) is nonlinear and Markovian in the unconditional master equation [62] in $\rho$. Under WCM, a Wiener process $dW$ with a Gaussian distribution [62] in the last term of Eq. (II.2) arises from the Gaussian-weighted projection over the eigenstates that allows the quantum information to be extracted continuously in the time domain, subject to stochastic disturbances in the last term of Eq. (II.2) and quantum decoherence in the penultimate term of Eq. (II.2). (Appendix B provides a detailed derivation of the SME.) The Lindblad operator $\mathcal{D}$ and the measurement superoperator $\mathcal{H}$ in Eq. (II.2) are given by
$\mathcal{D}(\hat{A})\rho$ $\equiv\hat{A}\rho\hat{A}^{\dagger}-\frac{1}{2}(\hat{A}^{\dagger}\hat{A}\rho+\rho\hat{A}^{\dagger}\hat{A}),$ (5)
$\mathcal{H}(\hat{A})\rho$ $\equiv\hat{A}\rho+\rho\hat{A}^{\dagger}-\langle\hat{A}+\hat{A}^{\dagger}\rangle\rho,$ (6)
with $\langle\hat{A}\rangle\equiv\textnormal{Tr}[\hat{A}\rho]$. The two operators serve to weakly drive the quantum state into the corresponding eigenstates to some degree. The number $n_{traj}$ of trajectories can be selected according to the following considerations. If the observable is the expected quantity, e.g., $\langle\hat{A}\rangle\equiv\textnormal{Tr}[\hat{A}\rho]$ with the observable $\hat{A}$ over the quantum state, using one trajectory is sufficient for the quantum information associated with the state to be extracted: $n_{traj}=1$. Experimentally, however, WCMs are often performed, e.g., of the current, which encodes the Wiener process in the observation and results in a large variance from the expectation value. To reduce the variance, more quantum trajectories should be used. To make computations feasible, we use five trajectories: $n_{traj}=5$.
A general quantity to measure the entanglement between arbitrary quantum bipartite systems for any mixed states is the logarithmic negativity [74, 75, 76], without the influence of the vacuum bath [77]. In contrast, the conventional pure-state entanglement measures, such as the von Neumann and Rényi entropy, capture both quantum and classical correlations. Since the goal of our study is harnessing the entanglement between the cavity and oscillator modes, we focus on the logarithmic negativity:
$E_{N}(\rho)\equiv\log_{2}||\rho^{T_{i}}||_{1},$ (7)
where $||X||_{1}=\textnormal{Tr}\sqrt{X^{\dagger}X}$ is the trace norm of the partial transpose $\rho^{T_{i}}$ with respect to the two subsystems $i=0$ (quantum-optical cavity mode) and $1$ (mechanical oscillator mode). The logarithmic negativity measures the degree to which $\rho^{T_{i}}$ fails to be positive, i.e., the extent of inseparability or entanglement, and it is the upper bound of the distillable entanglement [74, 75]. The logarithmic negativity is the full entanglement monotone [75], which satisfies the following criteria [78, 76]: (1) $E_{N}$ is a non-negative functional, (2) $E_{N}$ vanishes if the state $\rho$ is separable, and (3) $E_{N}$ does not increase on average under Gaussian local operations and classical communication [79, 80] or positive partial transpose preserving operations [81]. Since $E_{N}$ quantifies the quantum correlation between the bipartite systems in spite of the coupling to the vacuum bath, the value of $E_{N}$ calculated from the cavity mode is equal to that of the oscillator mode: $E^{0}_{N}=E^{1}_{N}=E_{N}$, which can be verified numerically.
In our computations, all the control actions $G$, $\alpha_{L}$, or detuning $\Delta$, and the dissipation coefficients $(\kappa,\gamma)$ are in units of $\omega_{m}$. The time unit is $\omega_{m}^{-1}$.
III Deep reinforcement-learning control of quantum optomechanical systems with linear interactions

### v2

@@ Abstract
Entanglement is fundamental to quantum information science and technology, yet controlling and manipulating entanglement — so-called entanglement engineering — for arbitrary quantum systems remains a formidable challenge. There are two difficulties: the fragility of quantum entanglement and its experimental characterization. We develop a model-free deep reinforcement-learning (RL) approach to entanglement engineering, in which feedback control together with weak continuous measurement and partial state observation is exploited to generate and maintain desired entanglement. We employ quantum optomechanical systems with linear or nonlinear photon-phonon interactions to demonstrate the workings of our machine-learning-based entanglement engineering protocol. In particular, the RL agent sequentially interacts with one or multiple parallel quantum optomechanical environments, collects trajectories, and updates the policy to maximize the accumulated reward to create and stabilize quantum entanglement over an arbitrary amount of time. The machine-learning-based model-free control principle is applicable to the entanglement engineering of experimental quantum systems in general.
Introduction
@@
Entanglement [1, 2, 3, 4, 5] is fundamental to all fields in quantum information science such as quantum sensing [6], quantum computation [7], and quantum networks [8, 9, 10, 11, 12, 13]. However, the inherent fragility of quantum entanglement and coherence [14] poses significant challenges for experimental applications. For example, in quantum computing, the application of quantum gates to quantum states needs to last for a finite amount of time [15, 16, 17, 18, 19, 20], making it critical to maintain the entanglement after its creation. Moreover, the transition from noisy intermediate-scale systems [21] to large-scale, fault-tolerant systems [16] requires sophisticated entanglement engineering strategies to establish and maintain entanglement through optimal control protocols in the presence of noise and decoherence.
At the present, a major limitation/challenge in entanglement engineering is the experimental observation design. Existing machine-learning based works use the full fidelity, i.e., the overlap between the current and target quantum states, as the observation metric. Applications range from the generation of two [22] and multi-qubit entangled states [23, 24] to specific many-body states [25, 26, 27] and single-particle quantum state engineering via deep reinforcement learning (RL) [28, 29]. However, full fidelity observation is not universally applicable in experiments. Moreover, obtaining the relationship between the entanglement and experimental observables is difficult. So far there have been no systematical methods to extract quantitative entanglement from experimental observation for arbitrary quantum systems [30, 31, 32], in spite of some initial exploration for specific systems. For example, an entanglement criterion for non-Gaussian states in coupled harmonic oscillators was developed [30]. Under the strong laser approximation, a Bell inequality was tested with photon counting [31], and stationary entanglement for Gaussian states was inferred from the continuous measurement of light only [32].
In this paper, using quantum optomechanical systems with linear or nonlinear photon-phonon interactions as a paradigm, we develop a deep RL approach to entanglement engineering. For quantum control of optomechanical systems, most existing theoretical studies focused on Gaussian states or the linear interaction regime [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 38, 44, 45], with the primary goal of generating entanglement as quickly as possible (entanglement enhancement) [43, 38, 44, 45]. Previous control methods are mostly model-based: prior information about the system model is needed, such as the pulse method [33, 34, 35, 36, 37], time-continuous laser-driven approaches [38, 39], periodic modulations [40, 41, 42], optimal pulse protocols [43], linear quadratic-Gaussian (LQG) methods [38], and coherent feedback methods using auxiliary optical components [44, 45]. We note that there were two previous works [46, 47] on model-free RL for controlling and stabilizing a quantum system with an inverted harmonic potential and a double-well nonlinear potential, respectively, to a target state using weak-current measurements (WCMs) and partial state observation. However, these two works did not address entanglement control, while our work is developing a model-free deep-RL method to realize non-Gaussian entanglement engineering using only photon number counting from WCMs. To our knowledge, prior to our work, model-free deep RL feedback control to create and stabilize the entanglement with WCM observations had not been available.
The particular aspects of our work that go beyond the existing works are briefly described, as follows. In our work, in the linear (nonlinear) interaction regime, the observation is the WCM photocurrent (the expectation value of the photon number). We note a previous work [29] that employed a proximal policy optimization (PPO) [48] RL agent, to generate different Fock states and the superposition of a single cavity mode based on observing the density matrix and a fidelity-based reward function. In contrast, the observable in our work is the photocurrent that is more experimentally accessible [49]. For quantum measurement, we use WCM in real-time feedback control, taking into consideration the resulting quantum stochastic process [50, 51, 47], and identify a numerical relationship between the entanglement and photocurrent. In both the linear and nonlinear regimes, we focus on non-Gaussian state control because, according to the nonlinear quantum master equation resulting from WCM, the time evolving quantum states are intrinsically non-Gaussian. Our deep RL control scheme is model-free [52], where policies or value functions are directly learned from the interactions with the quantum environment without any explicit model of this environment. This should be contrasted to the model-based deep RL methods [53], where a pre-built model of the environment for policy decision-making is needed. We demonstrate that, under the actions of the well-trained PPO or recurrent PPO RL agent, entanglement between the quantum optical and mechanical modes can be created and maintained about the target entanglement.
@@
Our main results are as follows. First, under the strong laser approximation, the interaction resulting from the radiation pressure between the cavity and the mechanical oscillator modes can be linearized and described by the beam-splitter Hamiltonian. During the training phase, the PPO agent interacts with parallel quantum environments and collects the subsequent data by episodic learning, with the observation being the WCM photocurrent. The deep-RL method can extract useful information from the measurement photocurrent, which is encoded in the Wiener process, and achieve the target entanglement engineering in a model-free manner for the quantum system that is dissipative due to coupling to the vacuum bath and is driven by a laser. In the testing phase, with the agent interacting and observing a single quantum environment, we demonstrate that the entanglement-engineering performance of our deep-RL method with WCM observation greatly exceeds that of both state-based Bayesian methods [54, 47] and random control. Second, when the driving laser field is not strong, the quantum optomechanical interaction is nonlinear [55, 56]. In this case, we articulate two training phases for nonlinear entanglement engineering. The first phase is utilized to infer the entanglement by the model-free deep RL, dubbed as the target-generating phase, where the observation of the PPO agent [with multilayer perceptions (MLPs)] is the logarithmic negativity and the reward function is constructed to limit the high-level excitation and facilitate entanglement learning. (Direct experimental measurement of the logarithmic negativity is currently not available.) The time series of the expected photon number in the regime from converged training episodes is selected as the target for the next phase. The second phase is then the target-utilization phase, where the recurrent PPO (with long short-term memory (LSTM) [57] added after MLPs) observes the expected photon number and obtains the reward only based on the target expected photon number obtained from the last phase. In this framework, the recurrent PPO controls the quantum state in the low-energy regime with the desired entanglement created and stabilized.